SOLVATION OF HCl, AGGREGATION OF H_{2}O, HCl AND THEIR DEUTERATED ISOTOPES IN SUPER-FLUID He DROPLETS^{a}

\underline{GERHARD W. SCHWAAB}, OZGÜR BIRER, ANNA GUTBERLET, MARKUS ORTLIEB, MAR-TINA HAVENITH, Department of Physical Chemistry II, Ruhr-University Bochum, Universitätsstr. 150, D-44780.

We investigate aggregates of HCl, H_{2}O and their deuterated isotopes embedded in superfluid helium using a helium nanodroplet spectrometer combined with a cw infrared OPO set up with an output power of up to 2.7 W and a resolution of better than 0.001 cm^{−1}. Depletion spectra of the super cooled aggregates (0.37 K) were recorded between 2650 cm^{−1} and 2720 cm^{−1}. We were able to assign the HCl stretch vibration of the HCl-H_{2}O complex. We also observed spectral features of (HCl)_{m}(H_{2}O)_{n} aggregates and isotopes in the region of the H_{3}O^{+} stretch vibration. The question of how many H_{2}O molecules are needed to solvate HCl will be addressed.

athis work was supported by FOR 618 of the German Science Foundation